OpenStudy (anonymous):

Find a polynomial of minimum degree (there are many) that has the given zeros 1- sq. of 3 and 1+ sq. rt. of 3

hartnn (hartnn):

if a is a zero, x-a is a factor. so here factors are (x-1+\(\sqrt3\))(x-1-\(\sqrt3\)) =\((x-1)^2\)-\((\sqrt3)^2\) =x^2-2x+1-3 =x^2-2x-2.

OpenStudy (anonymous):

oh my god youre not even explaining it but still i can understand it oh em gee how old are you

hartnn (hartnn):

i m sorry, do u need step wise explanation ?

OpenStudy (anonymous):

no i get it now

hartnn (hartnn):

ok, ask if there are any doubts in any step.